A stock has an expected return of 10.6 percent, its beta is 0.99, and the risk-free rate is 6.2 percent.
Required:

What must the expected return on the market be? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).)

  Expected return %
Answer #1

expected return=risk-free rate +Beta*(market rate- risk-free rate )

10.6=6.2+0.99*(Market rate-6.2)

(10.6-6.2)/0.99=Market rate-6.2

Market rate=(10.6-6.2)/0.99+6.2

=10.64%(Approx).